When 676J of heat energy is added to a 47G sample of solid copper the temperature increases to 88.6 Celsius what was the original temperature of the solid in Celsius the specific heat capacity of solid copper is 0.385 J/G•K


1
Expert's answer
2021-03-08T06:11:48-0500

H=mcΔt

Δt = H/mc

=676/(47 × 0.385)

= 37.36

Δt=t2-t1

t1 = t2-Δt

= 88.6 - 37.36 = 51.24°C